Complete the following table for a rectangle
perimeter 80 cm length 28 cm how cm of breadth
Send fomula of breadth

Answer by FrankM(1040)

perimeter 80 cm
length 28 cm
breadth = ?
We typically call the dimensions length and width. Here, breadth is used instead of width.
Perimeter is simply the sum of the 4 sides, i.e. 2L+2B
2L+2B=80
2(28)+2B=80
56+2B=80
2B=24
B=12cm